OpenStudy (anonymous):

The range of a set of data is 75. If the lowest value in the data set is 5, what is the highest value?

OpenStudy (anonymous):

5+75=80

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

range is max-min=80-5=75

OpenStudy (anonymous):

thank u :D i swear i learn bette here then in school

OpenStudy (anonymous):

naturally, excellent tutors here!

OpenStudy (anonymous):

yup yup!